A body is moving ina straight line  starting from rest moving with acceleration X for some time and then with uniform velocity V for some time and then regards with retardation B and comes to rest.total distance traveled is L then total time of journey is

Grade:7

2 Answers

Arun
3 years ago
Dear student
 
There is no image of question. Please check and repost the question with the figure. I will be happy to help you.
Vikas TU
3 years ago
Dear student 
T = t1 + t2 
v = u+at 
v = 0+ At1............1 
Decelerating 
0=V -Bt2........2
At1 = Bt2 
At1/B = t2 
T = t1 + At1/B